### v1

# Automated Retinal Layer and Fluid Segmentation and Cross-sectional Analysis using Spectral Domain Optical Coherence Tomography Images for Diabetic Retinopathy

Abstract
Background: Diabetic retinopathy (DR) is a leading microvascular complication of diabetes and a significant cause of vision loss. Without timely intervention, it can lead to irreversible blindness. Spectral Domain Optical Coherence Tomography (SD-OCT) enables high-resolution imaging of retinal structures, aiding in the early detection of pathological changes. AI-driven segmentation enhances diagnostic precision and reduces clinician workload. This study aims to develop a deep learning-based pipeline for automated retinal segmentation and thickness analysis to improve DR assessment and monitoring.
Methods: A deep neural network (DNN) was trained for automated segmentation of ten retinal layers, intra-retinal fluid, and hyperreflective foci (HRF). Initial manual annotations from five SD-OCT volumes were used to train the model, with iterative refinements improving accuracy. Performance was evaluated using five-fold cross-validation across four DNN architectures. Retinal thickness was measured via the K-nearest neighbours (KNN) algorithm, and Early Treatment Diabetic Retinopathy Study (ETDRS) maps visualized spatial variations. Statistical comparisons were made between Non-Proliferative (NPDR) and Proliferative Diabetic Retinopathy (PDR). The correlation between layer thickness and visual acuity within each DR group is also evaluated.
Results: SwinUNETR achieved the highest segmentation accuracy, with a Dice Similarity Coefficient (DSC) of 0.7719 and a Normalized Surface Dice (NSD) of 0.8149. VM-Unet showed superior DSC performance in specific layers (Vitreous, RNFL, GCL+IPL, ONL+IS), while SwinUNETR excelled in EZ, OS, RPE, and HRF regions. PDR patients exhibited increased thickness in the inner nasal and temporal of RNFL, whereas NPDR showed greater thickness in specific sectors of GCL+IPL, OS, and HRF sectors. Several sectors of INL and ONL+IS in NPDR show a significant correlation between layer thickening and better vision, but the reverse is true in PDR for RNFL, OPL, ONL+IS, and EZ. The accumulation of both fluid and HRF is also significantly correlated with vision decline in PDR, but no significance is found in NPDR. Similar findings can be derived from the predicted segmentation.
Conclusion: This study introduces an AI-based pipeline for automated DR assessment. It reduces manual annotation effort while maintaining high accuracy. The generated thickness maps provide valuable clinical insights, supporting disease monitoring and treatment planning.
Index Terms:
Optical Coherence Tomography; Diabetic Retinopathy; Layer and fluid segmentation; Retinal thickness Analysis;
I Introduction
Diabetic retinopathy (DR) is a prevalent microvascular complication of diabetes mellitus (DM) and a leading cause of vision impairment worldwide (Klein, 2007). It is estimated that approximately 20% of diabetic individuals over the age of 50 will develop DR, which, if left untreated, can progress to severe visual impairment or blindness (Teo et al., 2021). The progression of DR is influenced by several risk factors, including prolonged diabetes duration, elevated glycated hemoglobin (HbA1c) levels, hypertension, hyperlipidemia, obesity, and smoking (Klein et al., 2014)(Zhou et al., 2025)(Mori et al., 2024)(Cheung et al., 2016)(Zhang et al., 2024)(Cao et al., 2017). Clinically, DR is categorized into two primary stages: Non-Proliferative Diabetic Retinopathy (NPDR) and Proliferative Diabetic Retinopathy (PDR). NPDR represents an early stage, often asymptomatic, characterized by microvascular abnormalities that progressively compromise retinal capillary integrity. Without timely medical intervention, NPDR can advance to PDR, a more severe stage marked by pathological neovascularization due to chronic retinal ischemia. This progression increases the risk of severe complications such as vitreous hemorrhage and retinal detachment, ultimately threatening vision.
Spectral Domain Optical Coherence Tomography (SD-OCT) is a cutting-edge, non-invasive imaging technique that provides high-resolution, cross-sectional visualization of retinal structures. Its real-time image acquisition capability makes it an invaluable tool for DR screening and early diagnosis (Gabriele E. Lang, 2007). SD-OCT enables the detection of subclinical retinal changes by quantifying variations in retinal thickness and identifying fluid accumulation. Accurate and reliable segmentation of retinal layers and pathological features is crucial for DR diagnosis and treatment planning. Figure 1 illustrates a representative SD-OCT B-scan with retinal layer and fluid segmentation. The segmentation delineates nine essential retinal layers: the Retinal Nerve Fiber Layer (RNFL), Ganglion Cell Layer and Inner Plexiform Layer (GCL+IPL), Inner Nuclear Layer (INL), Outer Plexiform Layer (OPL), Outer Nuclear Layer and Inner Segment Layer (ONL+IS), Ellipsoid Zone (EZ), Outer Segment Layer (OS), and Retinal Pigment Epithelium (RPE). The region above the Internal Limiting Membrane (ILM) is also identified as the Vitreous, while the Choroid lies beneath Bruch’s Membrane (BM). Fluid segmentation involves three primary fluid types: intraretinal fluid (IRF), subretinal fluid (SRF), and pigment epithelial detachment (PED), all of which appear as hypo-reflective regions between the ILM and BM. Furthermore, hyperreflective foci (HRF), which manifest as high-intensity dot-like or clustered lesions, are also segmented.
Figure 1: Example of retinal OCT segmentation(image on the right) for a single B-scan(image on the left). This image is acquired from a 59-year-old female patient with NPDR. Ten retinal layers are segmented from top to bottom, plus the fluid and HRF within the retinal body.
Numerous studies have explored automated segmentation techniques for retinal layer analysis. Herzog et al. proposed an edge maximization and smoothness-constrained thresholding approach to delineate retinal boundaries (Herzog et al., 2004). Chiu et al. utilized a graphical cut algorithm to minimize the weighted sum of edge paths along connected nodes, effectively segmenting retinal layers (Chiu et al., 2010). Wang et al. introduced a multi-step approach that includes artifact removal, contrast enhancement, and segmentation via level set methods, k-means clustering, and Markov random fields (MRFs) (Wang et al., 2015). Traditional machine-learning techniques have also been employed for fluid segmentation. González et al. identified dark fluid regions in OCT scans using support vector machines (SVM) and random forest classifiers (Gonzalez et al., 2013). Chen et al. applied a graph-cut classifier followed by a region-growing algorithm for cystoid macular edema (CME) segmentation (Xinjian Chen et al., 2012). However, these conventional approaches are limited by their reliance on handcrafted features and their susceptibility to performance degradation in severely diseased cases.
Deep learning has emerged as a powerful alternative for automated retinal segmentation, offering greater robustness against variations in image quality and pathological abnormalities. Liu et al. utilized a ResNet-based convolutional neural network (CNN) combined with a random forest classifier for patch-wise layer segmentation (Liu et al., 2019). Kugelman et al. proposed a recurrent neural network (RNN) with a graph search framework to segment retinal layers in both healthy individuals and patients with age-related macular degeneration (AMD) (Kugelman et al., 2018). Hu et al. developed a multi-scale CNN capable of capturing different feature levels for improved segmentation accuracy (Hu et al., 2019). U-Net and its derivatives have become widely adopted among deep-learning models for medical image segmentation. U-Net’s encoder-decoder architecture, enhanced by skip connections, enables efficient spatial information preservation and mitigates vanishing gradient issues (Ronneberger et al., 2015). It has been successfully applied to retinal layer segmentation, fluid detection, and HRF analysis, achieving state-of-the-art (SOTA) performance (Roy et al., 2017)(Ma et al., 2021)(Tennakoon et al., 2018)(Schlegl et al., 2018). More recently, Vision Transformers (ViTs) have outperformed CNNs in large-scale datasets. Unlike CNNs, which rely on local receptive fields, ViTs employ self-attention mechanisms to capture global dependencies, which is particularly beneficial for detecting diffuse fluid regions. Xue et al. implemented a Swin-Transformer-based architecture for fluid segmentation in diabetic macular edema (DME) and AMD, demonstrating superior performance over traditional CNN-based models (Xue and Du, 2024). Kulyabin et al. leveraged the Segment Anything Model (SAM) for retinal fluid segmentation, incorporating point and bounding box prompts to outperform U-Net in macular hole and fluid segmentation tasks (Kulyabin et al., 2024). Despite these advancements, most existing studies focus on either the retinal layer or fluid segmentation, with limited efforts dedicated to integrating segmentation outcomes with clinical statistical analysis.
Several studies have examined the relationship between retinal layer thickness, fluid accumulation, and DR severity. Browning et al. analyzed macular thickness across different DR severity levels and observed a correlation between macular thickening and increased risk of subclinical edema (Browning et al., 2008). Kim et al. investigated choroidal thickness alterations in DR and DME patients, reporting a significant increase in choroidal thickness as DR severity progressed from mild/moderate NPDR to PDR (Kim et al., 2013). Cho et al. assessed macular and peripapillary retinal thickness in DR subjects, identifying statistically significant differences in retinal thickness across seven anatomical regions between DR and control groups (Cho et al., 2010). Santos et al. demonstrated that fluid accumulation within the outer segment (OS) layer was significantly associated with central retinal thickness and visual impairment in DME patients (Santos T et al., 2024). These findings suggest that retinal layer thickness and fluid distribution are reliable biomarkers for DR diagnosis and progression monitoring.
This study introduces a fully automated pipeline integrating retinal layer and fluid segmentation with a statistical analysis of structural changes in DR patients. The key contributions include:
  1. 1.
    Development of a scalable pipeline for large-scale manual segmentation refinement.
  2. 2.
    Quantitative evaluation of multiple SOTA deep learning models for retinal layer and fluid segmentation.
  3. 3.
    Cross-sectional analysis of retinal structural changes and demographic variations, providing clinically relevant insights into DR progression.
II Methods
II-A Data Acquisition
116 SD-OCT volumes were acquired from Sankara Nethralaya Eye Care Hospital in India. The imaging data was obtained using the Cirrus HD-OCT 5000 (Carl Zeiss Meditec, Dublin, CA, USA). Seventeen OCT volumes were captured in Macular Cube mode with a 512 × 128 pixels resolution, covering a 6 × 6 mm² macular region centered on the fovea. The remaining 99 volumes were scanned in OCTA mode at 350 × 350 pixels, covering a 6 × 6 mm² macular region. Despite differences in scanning speed and resolution, Wong et al. reported no significant variation in macular thickness measurements between the two modes (Wong et al., 2024). Table I presents the demographic details of the two DR groups, showing no significant differences in age, diabetes duration, or visual acuity (p $>$ 0.05). However, the gender distribution differs due to the limited number of female patients in the PDR group.
Group N Age (Mean ± SD) Gender Duration of Diabetes (yrs) Visual Acuity (LogMAR)
NPDR 66 60.00 ± 8.85 Female: 34, Male: 32 15.94 ± 7.97 0.35 ± 0.35
PDR 50 56.84 ± 8.24 Female: 13, Male: 37 14.66 ± 8.60 0.47 ± 0.41
p-value - 0.05023 0.00457 0.41508 0.10351
TABLE I: Demographic information of experimental DR groups. Numerical values are presented as mean ± standard deviation (SD). The visual acuity is expressed in the logarithm of the Minimum Angle of Resolution (LogMAR). The p-values are calculated using the Student’s t-test.
II-B Pre-processing
To prepare the raw OCT volumes for further analysis, we perform several pre-processing steps:
  • The approximate retinal center in each B-scan is adjusted to align with the center along the axial direction. The axial retinal center is estimated by computing the average axial position of pixels whose axial intensity values are more significant than the lowest 20th percentile. This helps initialize a starting point for axial motion correction.
  • A 3D Bounded Variation (BV) smoothing technique is applied to suppress noise while preserving sharp structural boundaries. This provides a better contrast between manual labelling and correction.
  • Motion artifacts are corrected in each B-scan’s axial and lateral directions. Axial translations are determined through cross-registration using the moving average of the central B-scan as a reference. Lateral corrections are achieved by performing registration based on the adjacent B-scans’ discrete Fourier transform (DFT). Rotational adjustments are computed by transforming translations into polar space, using the moving average of the central B-scan as a reference.
II-C Manual Segmentation
Figure 2 illustrates that the segmentation process follows a structured workflow. Initially, five volumes are manually annotated from scratch. To enhance efficiency, manual segmentation is performed on every fifth B-scan, while the intermediate B-scans are interpolated under the assumption that adjacent B-scans share structural similarities. However, B-scans that exhibit significant structural changes are individually labelled. These five manually labelled volumes serve as the initial training dataset for a deep neural network (DNN), which is trained using a data split ratio of 3:1:1 for training, validation, and testing. The trained model is then used to generate segmentation predictions for an additional 20 volumes, which are subsequently reviewed and manually corrected. This iterative process continues, with the network being retrained on an expanded dataset each time, ensuring that all volumes undergo accurate segmentation and manual verification.
Figure 2: Manual segmentation pipeline. Multiple iterations were performed between DNN training and manual corrections. A thickness analysis was conducted after segmentation had been completed and verified.
II-D Networks
We investigate the performance of four deep neural network (DNN) architectures, which are either widely used for medical image segmentation or have demonstrated state-of-the-art (SOTA) performance in related tasks:
  • U-Net: A well-established medical image segmentation model that employs an encoder-decoder architecture with skip connections (Ronneberger et al., 2015). While effective, it may struggle with high-resolution inputs due to a lack of global contextual awareness.
  • SegFormer: Designed by Xie et al., this framework utilizes a hierarchical transformer encoder combined with a lightweight MLP decoder to enhance feature extraction (Xie et al., 2021).
  • SwinUNETR: Proposed by Hatamizadeh et al., this architecture replaces the CNN-based encoder in U-Net with a Swin Transformer encoder, enabling multi-scale feature extraction through a shifted windowing mechanism (Hatamizadeh et al., 2022).
  • VM-UNet: Developed by Ruan et al., this model introduces a state-space model (SSM) and an asymmetric encoder-decoder structure, balancing computational efficiency while maintaining a global contextual view (Ruan et al., 2024).
The U-Net model is configured with a depth of five channels and three residual units. The 2D variant of SwinUNETR is used for training B-scans. The architectures of the remaining three networks are adopted as described in their original publications. The VM-UNet model is initialized using VMamba-S, pre-trained on the ImageNet-1k dataset (Liu et al., 2024).
II-E Training Settings
We employed 5-fold cross-validation, stratified by DR diagnosis, with a 4:1:1 ratio for training, validation, and testing. Each input consisted of a 3-channel image constructed by three repetitions of a single B-scan. Each training B-scan is resized to $512\times 512$. To mitigate class imbalance, excess Vitreous and Choroid regions were cropped. Various augmentation techniques are applied, including lateral flipping, Gaussian noise addition, contrast enhancement, rotation within the B-scan plane, and random intensity shifting.
For loss functions, we use the combinations of Dice loss, cross-entropy (CE) loss, and the L1 loss of texture differences. Given the ground truth label $y$ and predicted label $\hat{y}$, for every pixel $i$, the Dice loss is calculated as:
$L_{dice}(y,\hat{y})=1-\frac{2\cdot\sum_{i}y_{i}\hat{y_{i}}}{\sum_{i}y_{i}^{2}+\sum_{i}\hat{y_{i}}^{2}+\epsilon}$ (1)
We set $\epsilon$ to $10^{-6}$ to avoid the division by zero problem. The CE loss is defined as:
$L_{CE}(y,\hat{y})=-\frac{1}{N}\sum_{i}\left[y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})\right]$ (2)
The Sobel operator calculates gradients in horizontal($G_{x}$) and vertical($G_{y}$) directions, and the total gradient magnitude $G$ is the Euclidean norm. Given the label $Y$, the gradients are calculated as:
$G_{x}(Y)=Y\ast\begin{bmatrix}-1&0&1\\ -2&0&2\\ -1&0&1\end{bmatrix},\quad G_{y}(Y)=Y\ast\begin{bmatrix}-1&-2&-1\\ 0&0&0\\ 1&2&1\end{bmatrix}$
$G(Y)=\sqrt{G_{x}(Y)^{2}+G_{y}(Y)^{2}}$ (3)
The texture loss is defined as the L1 norm between the predicted and ground truth labels:
$L_{texure}(y,\hat{y})=\frac{1}{N}\sum_{i}\left|G_{i}(y)-G_{i}(\hat{y})\right|$ (4)
Thus, the total loss function is calculated as:
$L(y,\hat{y})=\alpha\cdot L_{dice}+\beta\cdot L_{CE}+\gamma\cdot L_{texure}$ (5)
The $\alpha$, $\beta$, and $\gamma$ are weighting factors for Dice, CE and texture losses, respectively. We empirically set $\alpha=\beta=\gamma=1$ for our experiment.
We empirically assign different class weights to CE loss to emphasize the class imbalance issue. Specifically, we assign 0.1 to Vitreous and Choroid, 0.5 to the rest of the layers, and 1 to fluid and HRF. We use AdamW optimizer with CosineAnnealing scheduler with the warm restart. We use the distributed parallel learning supported by the PyTorch Lightning module [^1] https://lightning.ai/, with a batch size of 8 and a learning rate of 1e-4. The training is deployed on NVIDIA V100 Volta GPU allocated by Cedar Compute Canada [^2] More information can be found at: https://docs.alliancecan.ca/wiki/Cedar.
II-F Evaluation Metrics
We evaluate the segmentation performance by overlapping areas and boundary alignment. We use the Dice similarity coefficient (DSC) to measure the similarity between the predicted and ground truth masks. Given correctly predicted pixels as True Positives(TP), incorrectly predicted pixels as False Positives(FP), and missing predicted pixels as False Negatives(FN), the Dice score is calculated as:
$Dice=\frac{2\cdot TP}{2\cdot TP+FP+FN}$ (6)
Nikolov et al. proposed the normalized surface Dice (NSD) to estimate the deviation of surface contours within a certain threshold $\tau$(Nikolov et al., 2018). Defining a set of Euclidean distances from predicted segmentation $\hat{Y}$ to ground truth segmentation $Y$ as $\mathcal{D}_{\hat{Y}Y}$, and vice versa as $\mathcal{D}_{Y\hat{Y}}$, we obtain the subset of distances that are smaller or equal to the threshold $\tau$ as:
$\mathcal{D}_{\hat{Y}Y}^{\prime}=\{d\in\mathcal{D}_{\hat{Y}Y}|d\leq\tau\}$
$\mathcal{D}_{Y\hat{Y}}^{\prime}=\{d\in\mathcal{D}_{Y\hat{Y}}|d\leq\tau\}$ (7)
The NSD is calculated as :
$NSD=\frac{|\mathcal{D}_{\hat{Y}Y}^{\prime}|+|\mathcal{D}_{Y\hat{Y}}^{\prime}|}{|\mathcal{D}_{\hat{Y}Y}|+|\mathcal{D}_{Y\hat{Y}}|}$ (8)
Special attention is needed for fluid evaluation. For True Negative(TN) cases where both ground truth and predicted fluid are absent, the NSD score should be the correct prediction. For False Positive(FP) and False Negative(FN) cases where the fluid is only present in one of the ground truths or predicted segmentations, the NSD score should be zero as the incorrect prediction. We set $\tau$ to $10$ pixels for all classes, roughly $3\%$ of the shortest image edge. The model performance will be evaluated without any of the post-processing steps mentioned in the original papers.
Additionally, we define the Under-Segmentation Score (USS) and the Over-Segmentation Score (OSS) to evaluate if the model fails to detect certain regions or assigns excessive labels to a class. Given the confusion matrix for N classes:
$CM=\begin{bmatrix}\text{TP}_{1}&\text{FP}_{1,2}&\text{FP}_{1,3}&\dots&\text{FP}_{1,N}\\ \text{FN}_{2,1}&\text{TP}_{2}&\text{FP}_{2,3}&\dots&\text{FP}_{2,N}\\ \text{FN}_{3,1}&\text{FN}_{3,2}&\text{TP}_{3}&\dots&\text{FP}_{3,N}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \text{FN}_{N,1}&\text{FN}_{N,2}&\text{FN}_{N,3}&\dots&\text{TP}_{N}\end{bmatrix}$ (9)
We compute the USS and OSS for certain class C as:
$USS_{C}=\frac{\sum_{\forall j\neq C}\text{CM}[C,j]}{\sum\text{CM}[C,:]}$ (10)
$OSS_{C}=\frac{\sum_{\forall i\neq C}\text{CM}[i,C]}{\sum\text{CM}[:,C]}$ (11)
A higher USS score indicates that a significant portion of the ground truth class $C$ was not detected, leading to under-segmentation, while a higher OSS score suggests that the model over-predicts class $C$, leading to over-segmentation.
II-G Thickness Analysis
Layer thickness computation is performed using the K-Nearest Neighbors (K-NN) algorithm. The layer boundaries are converted into 3D point clouds. For each data point on the upper layer, the closest corresponding point on the lower layer is identified based on Euclidean distance. Due to their unbounded nature on one side, the Vitreous and Choroid layers are excluded from these calculations. Given the anatomical complexity of the foveal pit, the central region thickness is excluded from the analysis to ensure more reliable and interpretable measurements.
The thickness of fluid and hyperreflective foci (HRF) regions is computed through direct accumulation, which is subsequently projected onto an en-face view for visualization. The Early Treatment Diabetic Retinopathy Study (ETDRS) grid is employed to assess thickness variations systematically across different macular regions. As depicted in Figure 3, this grid divides the macula into three concentric circles with diameters of 1mmmm\$\mathrm{m}$m, 3mmmm\$\mathrm{m}$m, and 6mmmm\$\mathrm{m}$m, all centered on the fovea. These circles define the central, inner, and outer subfields, subdivided into four quadrants: superior, inferior, nasal, and temporal.
Figure 3: ETDRS diagram for both left and right eyes. The diameters of the central circle, inner ring, and outer ring were 1 mm, 3 mm, and 6 mm, respectively. Nine subfields are numbered and named as follows: 1-CS(Central Subfield), 2-SI(Superior Inner), 3-NI(Nasal Inner), 4-II(Inferior Inner), 5-TI(Temporal Inner), 6-SO(Superior Outer), 7-NO(Nasal Outer), 8-IO(Inferior Outer), 9-TO(Temporal Outer).
III Results
III-A Segmentation
Table II presents the segmentation results for four models, with values averaged across five-fold cross-validation. Tables II(a) and II(b) separately report the DSC and NSD metrics. A repeated-measures analysis of variance (ANOVA) is conducted to compute the F-statistics and p-values for each retinal layer in both DSC and NSD. The ”Avg” column represents the average performance across all layers, with the highest score for each layer highlighted in bold.
SwinUNETR achieves the highest overall DSC and NSD among the evaluated models, demonstrating superior segmentation performance, particularly in the OPL, Choroid, and HRF regions. VM-UNet exhibits competitive performance, achieving the best DSC and NSD scores in the Vitreous, RNFL, and fluid regions. U-Net and SegFormer perform comparably, though U-Net slightly outperforms SegFormer in DSC across most layers, whereas SegFormer demonstrates marginally better NSD performance.
Notably, high F-values and small p-values are observed in the RNFL, GCL+IPL, INL, ONL+IS, fluid, and HRF layers, indicating significant variability in segmentation accuracy across models within these regions. These findings suggest that model predictions are less consistent in these layers, potentially due to structural complexity or segmentation challenges inherent to these regions.
Model Layer Dice Vitreous RNFL GCL+IPL INL OPL ONL+IS EZ OS RPE Choroid Fluid HRF Avg.
U-Net 0.9887 0.8723 0.8928 0.8180 0.7714 0.9151 0.7247 0.7331 0.8510 0.9800 0.2522 0.4075 0.7672
SegFormer 0.9897 0.8722 0.8932 0.8157 0.7680 0.9139 0.7210 0.7254 0.8403 0.9772 0.2122 0.3228 0.7543
SwinUNETR 0.9871 0.8713 0.8961 0.8259 0.7788 0.9148 0.7175 0.7267 0.8440 0.9804 0.2806 0.4402 0.7719
VM-UNet 0.9899 0.8740 0.8988 0.8269 0.7768 0.9185 0.7120 0.7173 0.8396 0.9798 0.2813 0.4211 0.7697
F-value 18.86 161.84 95.87 50.10 37.33 58.13 14.44 13.58 13.74 5.569 33.97 52.64 -
p-value 4.130×$\times 10^{-5}$ 2.504×$\times 10^{-10}$ 5.322×$\times 10^{-5}$ 2.166×$\times 10^{-7}$ 1.110×$\times 10^{-6}$ 9.368×$\times 10^{-8}$ 1.545×$\times 10^{-4}$ 2.072×$\times 10^{-4}$ 1.961×$\times 10^{-4}$ 9.023×$\times 10^{-3}$ 1.861×$\times 10^{-6}$ 1.640×$\times 10^{-7}$
(a) Mean Dice Similarity Coefficient (DSC) with repeated-measure ANOVA results
Model Layer NSD Vitreous RNFL GCL+IPL INL OPL ONL+IS EZ OS RPE Choroid Fluid HRF Avg.
U-Net 0.9540 0.8860 0.8408 0.8555 0.8381 0.8414 0.9338 0.9252 0.8786 0.9160 0.2847 0.5187 0.8061
SegFormer 0.9604 0.8888 0.8455 0.8580 0.8438 0.8434 0.9343 0.9231 0.8700 0.9181 0.2253 0.4105 0.7934
SwinUNETR 0.9576 0.8890 0.8499 0.8661 0.8493 0.8497 0.9313 0.9199 0.8726 0.9207 0.3186 0.5546 0.8149
VM-UNet 0.9628 0.8899 0.8484 0.8644 0.8425 0.8452 0.9327 0.9190 0.8586 0.9167 0.3221 0.5418 0.8120
F-value 27.65 36.34 31.55 22.64 11.30 4.054 19.06 12.36 3.485 6.568 19.62 44.15 -
p-value 5.640×$\times 10^{-6}$ 1.288×$\times 10^{-6}$ 2.776×$\times 10^{-6}$ 1.614×$\times 10^{-5}$ 4.901×$\times 10^{-4}$ 2.632×$\times 10^{-2}$ 3.915×$\times 10^{-5}$ 3.238×$\times 10^{-4}$ 4.142×$\times 10^{-2}$ 4.864×$\times 10^{-3}$ 3.372×$\times 10^{-5}$ 4.389×$\times 10^{-7}$
(b) Mean Normalized Surface Dice (NSD) with repeated-measure ANOVA results
TABLE II: Comparison of segmentation performance across four models. Dice and NSD scores are calculated by averaging over five validation folds. The ”Average” column represents the mean performance across all regions per model. The best score in each layer is highlighted in bold. The bottom rows show the F-value and p-value from repeated-measures ANOVA for statistical significance.
Figure 4 presents the mean difference in DSC and NSD between each model pair across retinal regions. For a model pair denoted as ”A vs B,” the mean difference is computed by subtracting model A’s mean performance from model B’s. The significance of these differences is indicated by an asterisk marker preceding the model pair: ”***” denotes strong statistical significance ($p<0.001$), ”**” represents moderate significance ($p<0.01$), and ”*” indicates statistical significance at $p<0.05$. The absence of a marker signifies no significant difference.
Significant mean differences with strong statistical significance are observed in the fluid and HRF regions. SwinUNETR and VM-UNet consistently demonstrate superior performance across retinal layers. While U-Net and SegFormer perform competitively in most layers, their performance in fluid and HRF regions is inferior to that of SwinUNETR and VM-UNet.
SwinUNETR and VM-UNet exhibit comparable performance in the INL, OPL, Choroid, and fluid regions for DSC. However, VM-UNet significantly outperforms SwinUNETR in the Vitreous, RNFL, GCL+IPL, and ONL+IS layers, whereas SwinUNETR achieves considerably better performance than VM-UNet in the EZ, OS, RPE, and HRF regions. Regarding NSD, SwinUNETR demonstrates a significantly larger NSD than VM-UNet in the Vitreous, OPL, ONL+IS, RPE, Choroid, and HRF layers, with no significant differences observed in the remaining regions.
Figure 4: Comparison of the mean difference of DSC and NSD scores between models for each retinal region. For each pair of models shown in the subplot, the mean difference is calculated by subtracting the mean of the first model from the second model. The marker before the name of each model pair indicates different significant levels(p-values) calculated by the Tukey HSD method. ***: $p<0.001$, **: $p<0.01$, *: $p<0.05$, no marker: $p\geq 0.05$. Some barplots are invisible due to the limited precision of the mean difference.
Figure 5 presents examples of segmentation model predictions. Each sub-image displays five B-scans selected from the 60 central B-scans. The predicted segmentation is generated using the sub-fold model corresponding to the test set to which the patient belongs.
Figure 5(a) illustrates an NPDR patient exhibiting severe intraretinal fluid. All four models successfully segment the majority of fluid regions. However, U-Net and SegFormer demonstrate weaker fine-layer segmentation performance than SwinUNETR and VM-UNet, particularly in the RNFL and OPL layers. VM-UNet excels in preserving layer continuity and structural integrity, whereas U-Net and SwinUNETR exhibit discontinuities in the OPL layer in the fourth and fifth B-scans. Notably, VM-UNet tends to under-segment the fluid relative to the other models.
Figure 5(b) displays an NPDR patient with pronounced HRF. Shading artifacts beneath large HRF clusters disrupt the continuity of the lower layers. VM-UNet demonstrates superior performance in maintaining layer integrity despite losing pixel intensity in the bottom three B-scans. In contrast, U-Net and SegFormer struggle to compensate for these artifacts, while SwinUNETR erroneously misclassifies portions of HRF within the Choroidal region.
Figure 5(c) depicts a PDR patient with severe fluid accumulation. U-Net exhibits the weakest performance in fluid segmentation among all models, particularly in the third and fourth B-scans. Additionally, all models show varying degrees of under-segmentation in the final B-scan.
Figure 5(d) illustrates a PDR patient with partial retinal detachment. VM-UNet demonstrates robust performance across all Example B-scans, providing a more accurate segmentation of the RNFL compared to the other models.
(a) NPDR patient with severe fluid accumulation.
(b) NPDR patient with HRF and shading artifacts.
(c) PDR patient with severe fluid accumulation.
(d) PDR patient with partial retinal detachment.
Figure 5: Comparison of OCT B-scan segmentation results across different retinal conditions. Each row represents a different B-scan, while columns correspond to different segmentation models and patient conditions.
Figure 6 shows the USS and OSS for the top-2 performance models SwinUNETR and VM-UNet. VM-UNet shows lower USS than SwinUNETR in most regions except for HRF, and it has less over-segmentation in most areas except for INL, OPL, EZ and OS. Overall, both models tend to under-segment in OPL, EZ and OS layers, plus the fluid and HRF. Over-segmentation is observed in the INL, OPL, ONL+IS, EZ and OS layers.
III-B Thickness
Statistical significance between the retinal thickness and DR groups is analyzed in Table III. The thickness analysis is applied to the predicted segmentation of the top-2 performance models SwinUNETR and VM-UNet for comparison with the ground-truth segmentation. For each layer sector, the p-values of Welch’s t-test (first row) and Analysis of Covariance (ANCOVA) adjusted by age, gender and diabetes duration (second row) are calculated. The p-values less than 0.05 are considered statistically significant. For ground-truth results, PDR has a significantly larger thickness than NPDR in the NI and II sectors of RNFL before and after p-value adjustment, but vice versa in II of GCL+IPL and all sectors of the OS layer except TI. NPDR has a significantly larger thickness in TI and IO of the GCL+IPL layer, TI of the OS layer, and more extensive volume accumulation in NI and TI of HRF. However, the significance is lost after adjusting for confounders. No significance is found in other layer sectors.
For model prediction results, SwinUNETR shows no significance in RNFL, while VM-Unet preserves the significance of ground-truth results but exaggerates them in the SI, SO, and TO sectors. SwinUNETR shows consistent significance in the IO sector of GCL+IPL, while VM-Unet only shows significance in II. Both models maintain the significance of the OS layer in most sectors except for NI and NO sectors but hyperbolize in the IO sector of RPE. No significance is shown in the fluid and HRF region for both models.
Fig 7 shows the thickness distribution across different retinal layers and sectors for NPDR and PDR groups. The outliers are removed outside the $5^{th}$ and $95^{th}$ percentiles, allowing a more robust interpretation of group differences. The diamond markers in each subplot show the mean thickness of each DR group without outlier removal. The PDR group has larger mean values and broader distributions in most sectors of RNFL, INL, OPL, EZ, and fluid, but it is vice versa for GCL+IPL, ONL+IS, OS, RPE and HRF.
Region Sector
SI NI II TI SO NO IO TO
Ground-truth
RNFL 0.152 0.045 0.110 0.011 0.094 0.053 0.448 0.537
0.183 0.033 0.083 0.012 0.133 0.052 0.550 0.623
GCL+IPL 0.201 0.364 0.025 0.050 0.865 0.536 0.045 0.246
0.230 0.447 0.035 0.084 0.764 0.604 0.060 0.377
INL 0.943 0.886 0.255 0.117 0.935 0.628 0.283 0.634
0.906 0.714 0.276 0.069 0.990 0.676 0.274 0.514
OPL 0.663 0.263 0.525 0.693 0.331 0.875 0.647 0.102
0.833 0.288 0.511 0.744 0.325 0.867 0.710 0.089
ONL+IS 0.314 0.713 0.101 0.218 0.109 0.910 0.105 0.630
0.333 0.588 0.072 0.228 0.101 0.724 0.070 0.674
EZ 0.722 0.759 0.562 0.884 0.921 0.498 0.737 0.468
0.649 0.652 0.503 0.702 0.998 0.496 0.807 0.500
OS 0.001 0.010 0.003 0.042 0.000 0.001 0.000 0.001
0.003 0.016 0.006 0.062 0.000 0.002 0.000 0.002
RPE 0.517 0.136 0.252 0.168 0.428 0.212 0.284 0.996
0.498 0.828 0.293 0.740 0.518 0.809 0.252 0.568
Fluid 0.517 0.136 0.252 0.168 0.428 0.212 0.284 0.996
0.517 0.101 0.213 0.159 0.519 0.283 0.381 0.825
HRF 0.535 0.634 0.033 0.250 0.032 0.114 0.322 0.572
0.681 0.528 0.059 0.284 0.072 0.120 0.384 0.580
SwinUNETR / VM-UNet
RNFL 0.183 / 0.039 0.094 / 0.027 0.349 / 0.142 0.504 / 0.017 0.076 / 0.028 0.149 / 0.089 0.594 / 0.777 0.059 / 0.007
0.201 / 0.039 0.063 / 0.014 0.259 / 0.107 0.518 / 0.020 0.110 / 0.042 0.125 / 0.076 0.607 / 0.861 0.062 / 0.007
GCL+IPL 0.387 / 0.436 0.461 / 0.408 0.053 / 0.018 0.135 / 0.158 0.921 / 0.726 0.413 / 0.296 0.044 / 0.068 0.463 / 0.435
0.401 / 0.478 0.543 / 0.507 0.075 / 0.024 0.213 / 0.263 0.918 / 0.679 0.397 / 0.240 0.048 / 0.080 0.603 / 0.611
INL 0.396 / 0.612 0.739 / 0.813 0.505 / 0.238 0.684 / 0.124 0.657 / 0.857 0.915 / 0.938 0.224 / 0.460 0.424 / 0.213
0.419 / 0.732 0.853 / 0.607 0.546 / 0.208 0.613 / 0.057 0.590 / 0.841 0.868 / 0.908 0.204 / 0.453 0.357 / 0.156
OPL 0.453 / 0.859 0.222 / 0.397 0.359 / 0.719 0.242 / 0.696 0.983 / 0.736 0.947 / 0.855 0.204 / 0.271 0.422 / 0.326
0.494 / 0.996 0.201 / 0.421 0.352 / 0.701 0.224 / 0.777 0.967 / 0.792 0.929 / 0.868 0.212 / 0.321 0.362 / 0.249
ONL+IS 0.154 / 0.468 0.729 / 0.937 0.117 / 0.364 0.087 / 0.128 0.083 / 0.150 0.712 / 0.982 0.023 / 0.066 0.427 / 0.497
0.166 / 0.495 0.564 / 0.967 0.086 / 0.298 0.110 / 0.155 0.061 / 0.120 0.462 / 0.838 0.014 / 0.053 0.482 / 0.578
EZ 0.329 / 0.339 0.564 / 0.466 0.330 / 0.153 0.179 / 0.230 0.558 / 0.216 0.279 / 0.681 0.549 / 0.312 0.668 / 0.908
0.474 / 0.464 0.667 / 0.622 0.409 / 0.252 0.368 / 0.437 0.613 / 0.231 0.229 / 0.629 0.538 / 0.344 0.649 / 0.974
OS 0.003 / 0.003 0.063 / 0.227 0.024 / 0.030 0.026 / 0.029 0.000 / 0.000 0.342 / 0.229 0.000 / 0.000 0.001 / 0.000
0.004 / 0.005 0.057 / 0.203 0.029 / 0.040 0.031 / 0.034 0.000 / 0.001 0.272 / 0.185 0.000 / 0.000 0.002 / 0.001
RPE 0.347 / 0.100 0.279 / 0.112 0.344 / 0.111 0.539 / 0.256 0.350 / 0.205 0.313 / 0.411 0.032 / 0.018 0.169 / 0.286
0.374 / 0.110 0.308 / 0.102 0.400 / 0.129 0.536 / 0.342 0.319 / 0.206 0.259 / 0.349 0.033 / 0.020 0.149 / 0.354
Fluid 0.277 / 0.599 0.076 / 0.108 0.183 / 0.353 0.108 / 0.158 0.916 / 0.128 0.910 / 0.384 0.125 / 0.497 0.486 / 0.416
0.277 / 0.618 0.053 / 0.083 0.159 / 0.322 0.100 / 0.154 0.864 / 0.165 0.983 / 0.418 0.067 / 0.532 0.386 / 0.345
HRF 0.716 / 0.845 0.564 / 0.681 0.061 / 0.125 0.523 / 0.694 0.064 / 0.068 0.353 / 0.113 0.467 / 0.279 0.871 / 0.788
0.880 / 0.973 0.432 / 0.560 0.102 / 0.171 0.524 / 0.714 0.120 / 0.129 0.343 / 0.101 0.553 / 0.367 0.835 / 0.776
TABLE III: Statistical significance of retinal thickness plus fluid and HRF volume differences for NPDR vs PDR. The first table shows results derived from the ground-truth label. The second table shows the results using the predicted segmentation, where each cell of two p-values represents the results of SwinUNETR and VM-UNet, respectively. Within each layer, the value on the first row represents the p-value from Welch’s t-test, and the second row represents the p-values adjusted by age, gender and duration of diabetes using Analysis of Covariance (ANCOVA). The p-values less than 0.05 are highlighted in bold.
Figure 9 presents four examples of thickness comparisons between NPDR and PDR groups using the ETDRS diagram described in Figure 3. Figures 8(a)9(b) correspond to cases of a 54-year-old male OD, a 57-year-old male OS, and a 59-year-old female OD and OS images, respectively. Each pair of patients is matched by age, gender, and eye laterality.
For each retinal region, the first row displays the En Face image overlaid with the corresponding layer thickness heatmap, while the second row presents the sector-wise quantitative average thickness. The En Face image is generated using each layer’s maximum intensity projection (MIP). For fluid and HRF, the En Face projection is derived from the entire retinal body (from the ILM to the BM), with thickness representing the accumulated volume in $\mu m^{3}$.
In the Example shown in Figure 8(a), PDR exhibits greater thickness than NPDR in nearly all RNFL sectors, except for the inner nasal and inferior regions. Similar trends are observed in the nasal and inferior OPL regions, inner sectors, and outer nasal sectors of fluid. Conversely, NPDR shows greater thickness in the inner sectors of the GCL+IPL, the inner superior INL, all OS sectors, and the temporal and superior HRF sectors.
In Figure 8(b), PDR demonstrates increased thickness in the nasal and superior RNFL sectors, the inner INL and OPL regions, nearly all fluid regions (except for the outer superior and temporal sectors), and all HRF regions. In contrast, NPDR exhibits greater thickness in the inner temporal and superior GCL+IPL regions and across all OS sectors.
For female patients shown in Figures 9(a) and 9(b), RNFL thinning is observed in the outer nasal and temporal sectors for PDR compared to NPDR. Additionally, GCL+IPL and RPE thinning are evident in the inner sectors of PDR. Fluid accumulation is significantly more pronounced in the inner sectors of PDR patients, whereas HRF is more prevalent in NPDR patients.
Figure 6: Comparison of USS(left) and OSS(right) for SwinUNETR and VM-UNet. Lower values indicate better segmentation performance.
Figure 7: The distribution of layer thickness and fluid volume measurements across different sectors for patients diagnosed with NPDR and PDR. The outliers are removed outside the $5^{th}$ and $95^{th}$ percentiles. The diamond markers represent the mean thickness for each group, calculated from the original data without outlier removal.
(a) Example of 54-year-old male patients with right eye image.
(b) Example of 57-year-old male patients with left eye image.
(a) Example of 59-year-old female patients with right eye image.
(b) Example of 59-year-old female patients with left eye image.
Figure 9: Examples of ETDRS thickness map comparison between NPDR and PDR patient with matched age, gender and eyeside. For each region, the first row shows the thickness heatmap overlayed onto the layer En Face image. The second row shows the quantitative average thickness for each sector.
We also evaluated the association between retinal layer sector thickness and visual acuity, as shown in table IV. The evaluation is performed for NPDR and PDR separately. Each value represents the Pearson correlation coefficient with a p-value within a specific layer sector. The p-values are adjusted by age, gender and diabetes duration. For NPDR, significant negative correlations are found in the SI sector of INL plus the II and SO sectors of ONL+IS. It suggests that layer thickening in these regions is associated with better vision. However, in PDR, we observe significant positive correlations in the II, SO, NO and IO sectors of RNFL, SO sector of OPL, NO sector of ONL+IS, TO sector of EZ, multiple sectors for both fluid (II, TI, SO, NO, TO) and HRF (SI, SO, NO). This finding indicates that thickening these layers and accumulations of fluid and HRF may result in worse vision.
Discussion
The segmentation performance varied across models, highlighting differences in architectural strengths and their ability to segment specific retinal layers and fluid-related abnormalities. SwinUNETR and VM-UNet consistently achieved high DSC and NSD scores, indicating their robustness in handling complex retinal structures. SwinUNETR particularly excelled in segmenting the OPL and HRF layers, which may be attributed to its transformer-based architecture that effectively captures long-range dependencies. VM-UNet, on the other hand, performed better in segmenting the fluid regions, suggesting that its sequential nature enhances segmentation continuity, particularly in areas with less distinct boundaries.
The performance differences in DSC and NSD indicate that while both models perform well, their strengths lie in different layers. DSC reflects volumetric segmentation accuracy, whereas SwinUNETR demonstrated better performance in HRF, likely due to its ability to differentiate small scattered hyper-reflective lesions. However, NSD provides insights into surface delineation accuracy, where VM-UNet showed superior performance in fluid regions, maintaining smoother boundaries compared to other models. U-Net and SegFormer, though competitive in some layers, exhibited weaker performance in fine layer segmentation, particularly in RNFL and OPL, where structural continuity is essential for accurate disease characterization. Although the SwinUNETR slightly outperforms VM-UNet in several regions, VM-UNet has significantly lower computational complexity (O(N)) than SwinUNETR (O($N^{2}$), which is crucial for remote deployment in clinics with limited computational resources.
The segmentation of fluid and HRF remains a significant challenge across all models. Fluid regions exhibit substantial variability, with VM-UNet demonstrating better spatial continuity but often under-segmenting these regions. In contrast, SwinUNETR captures fluid regions more extensively but is prone to occasional over-segmentation. HRF segmentation presents an even more significant challenge due to the presence of small, widely distributed hyper-reflective regions. Both models tend to under-segment fluid and HRF, frequently misclassifying them into adjacent retinal layers such as OPL and ONL+IS. Moreover, SwinUNETR generally exhibits a greater tendency to under-segment retinal regions than VM-UNet. Although significant weight adjustments were applied to fluid and HRF regions, as described in Section II-E, additional strategies are needed to enhance model learning and improve segmentation performance in these complex regions.
Significant differences in retinal layer thickness between NPDR and PDR offer valuable insights into the progression of DR. Prior to adjusting for potential confounders, including age, gender, and duration of diabetes, RNFL was significantly thicker in the NI and TI sectors in PDR. This finding contrasts with previous reports indicating RNFL thinning in more advanced stages of NPDR and PDR (Ulhaq et al., 2022)(Bhaskaran et al., 2023). It suggests that transient swelling, likely induced by abnormal blood vessel growth, may occur during the early acute phase of PDR but subsequently diminishes as chronic edema and neuroinflammation lead to tissue loss (Garcia et al., 2013). After adjusting for confounders, statistical significance persists in these two RNFL sectors, suggesting that these structural changes alone may serve as potential biomarkers for the early detection of PDR. When comparing these findings to model predictions, VM-UNet demonstrates greater consistency with ground-truth results than SwinUNETR, particularly in the RNFL and OS sectors. The loss of statistical significance for HRF is likely due to under-segmentation by both models. This highlights the sensitivity of retinal thickness analysis to segmentation variations and underscores VM-UNet’s suitability for detecting significant differences in retinal thickness between NPDR and PDR.
GCL+IPL exhibited significant thinning in PDR in the II, TI, and IO sectors before adjustment, reflecting the apoptosis of ganglion cells due to potential neurodegeneration and inflammatory responses. However, after adjusting for confounders, the TI and IO sectors are no longer significant, suggesting that systemic factors play a role in some but not all aspects of GCL+IPL thinning in PDR. Similarly, the OS layer showed considerable thinning in PDR across all sectors before confounders adjustment, highlighting widespread photoreceptor degeneration and atrophy. After adjustment, thinning in the TI sector lost significance. This suggests that overall OS deterioration is a hallmark of DR progression and an early marker of photoreceptor damage in PDR.
While not statistically significant in the t-test analysis, fluid accumulation was more prevalent in PDR than NPDR based on example cases shown in Figure 9. This suggests that while variability in fluid distribution may limit statistical significance, extensive fluid in PDR remains clinically relevant. The increased fluid volume aligns with the pathological features of PDR, where extensive vascular leakages lead to blood-retina barrier breakdown and, eventually, fluid accumulation. Before adjustment, HRF was significantly more widespread in NPDR than PDR, particularly in the II and SO sectors, supporting the hypothesis that HRF is an early marker of inflammation. However, this does not necessarily signify reduced disease severity of PDR but rather a transition to neovascularization toward deeper retinal layers. The chronic vascular leakage may cause the HRF to disperse or form into larger hard exudates and hemorrhages with less visibility in OCT. After adjusting for confounders, these differences lost statistical significance. This suggests that while HRF may be an important early indicator of neuroinflammation, its distribution may be influenced by demographic and disease-specific factors.
Our findings reveal significant associations between retinal layer thickness and visual acuity within NPDR and PDR patients. In NPDR, the observed negative correlations between thickness in specific retinal sectors and visual acuity suggest that moderate thickening in specific layers may have a protective effect with latencies. Conversely, positive correlations across multiple layers in PDR indicate that further increased thickness without proper treatment will lead to vision decline. Specifically, in NPDR, negative correlations in the SI sector of the INL and II/SO sectors of the ONL+IS suggest that mild thickening in these regions may represent a compensatory or adaptive response to metabolic activity, potentially maintaining structural integrity and delaying visual deterioration.
For PDR, thickening in several RNFL sectors (II, SO, NO, IO) is associated with poorer vision. This aligns with previous findings that transient edema and axonal swelling may occur in the early stages of PDR, coinciding with significant visual impairment. Additionally, increased thickness in the SO sector of the OPL, NO sector of the ONL+IS, and TO sector of the EZ is strongly linked to worse vision, likely due to edema-induced synaptic dysfunction between photoreceptors and bipolar cells, as well as EZ disruption affecting photoreceptor integrity.
Furthermore, fluid and HRF accumulation strongly correlate with PDR vision loss but not NPDR. The increased presence of fluid (II, TI, SO, NO) and HRF (SI, SO, NO) suggests progressive neurodegeneration and inflammatory activity in PDR, reinforcing the need for early intervention during the NPDR stage to prevent irreversible retinal damage.
While these findings provide valuable insights into the structural and functional relationships in NPDR and PDR, they are inherently result-oriented and restricted to the specific cohort analyzed. Future research should incorporate larger, multi-center cohorts with diverse demographic and clinical characteristics to enhance their generalizability. Validations in the interventional clinical trials will also help determine their potential applicability in guiding personalized treatment strategies for DR.
IV Limitation and Future Work
Despite the strengths of this study, several limitations must be acknowledged.
First, manual labelling imperfections impact both model performance and thickness measurements. Retinal layer segmentation is inherently challenging due to subtle boundary variations and overlapping structures. In cases where excessive fluid penetrates the layer boundaries, some portions of the layer become invisible or physically diminished. More clinical expertise is needed to segment the extreme instances properly. HRF segmentation suffers from inconsistencies in ground-truth annotations, as small, widely distributed foci are challenging to delineate manually. Interestingly, in some cases, automated models provided more precise segmentations than the manual ground truth, particularly for HRF regions. Figure 10 shows that the model prediction outperforms the ground truth in HRF segmentation. This suggests that deep learning models could be leveraged for segmentation and refining human annotations. Additionally, using more pre-processing and post-processing techniques may help improve the performance, such as the pixel-wise relative positional map as an extra input and random forest classifier as a label refiner(Ma et al., 2021).
Second, additional model comparisons may be necessary to provide a more comprehensive evaluation of segmentation approaches. While SwinUNETR and VM-UNet demonstrated superior performance, other architectures excel in certain perspectives. For example, MedSAM enables universal medical image segmentation with zero-shot capabilities(Ma et al., 2023). The self-supervised few-shot semantic segmentation can be used for a limited number of labels(Ouyang et al., 2022). A broader comparison across multiple deep learning models could offer more insights into the trade-offs between performance, efficiency, and generalizability.
Third, this study’s cross-sectional nature limits its ability to track disease progression over time. Longitudinal studies would provide better insights into how retinal layer thickness evolves in DR. For example, this study observes RNFL thickening in PDR, but longitudinal research often reports RNFL thinning due to neurodegeneration. This suggests that the transient swelling observed in cross-sectional studies may eventually lead to structural atrophy, which can only be confirmed through long-term follow-up studies.
Fourth, while the sample size is sufficient to detect significant differences, it may limit the generalizability of the findings. A larger dataset encompassing a broader range of DR severities and treatment histories could provide more robust conclusions. Expanding the cohort to include more diverse patient populations would also improve the applicability of the findings across different clinical settings.
Fifth, the lack of a detailed NPDR severity grading system may limit the ability to distinguish early, intermediate, and severe NPDR stages. Different NPDR severities likely exhibit distinct retinal layer changes, and a more granular classification system could enhance the understanding of DR progression. Future studies should explore integrating severity-based stratification to assess how thickness variations differ across NPDR subtypes. To our best knowledge, most DR grading datasets with public access focus on fundus color images like Messidor [^3] https://www.adcis.net/en/third-party/messidor/ and DRTiD [^4] https://github.com/FDU-VTS/DRTiD.
Lastly, integrating multi-modal imaging techniques such as OCT angiography (OCTA) could provide additional insights into the vascular changes associated with DR. For example, Alam et al. discovered the difference of vascular complexity features between NPDR and PDR patients(Alam et al., 2021). Multiple OCT parameters are significantly correlated with DR severity(Laotaweerungsawat et al., 2020). Combining structural OCT findings with functional vascular imaging may improve disease characterization and facilitate more targeted therapeutic interventions.
Region Sector
SI NI II TI SO NO IO TO
Non-Proliferative Diabetic Retinopathy (NPDR)
RNFL 0.0093 (0.9424) 0.0329 (0.7983) 0.0857 (0.5042) 0.0270 (0.8338) -0.0083 (0.9488) 0.0532 (0.6791) 0.0223 (0.8623) -0.0325 (0.8005)
GCL+IPL -0.0612 (0.6335) 0.0450 (0.7265) 0.0225 (0.8608) -0.0434 (0.7357) -0.0862 (0.5019) -0.0023 (0.9860) -0.0785 (0.5406) 0.0342 (0.7904)
INL -0.3464 (0.0054) -0.1157 (0.3666) -0.0615 (0.6320) -0.0166 (0.8974) 0.0017 (0.9895) -0.0479 (0.7094) -0.1082 (0.3986) -0.0842 (0.5116)
OPL -0.0642 (0.6171) -0.0515 (0.6883) -0.0777 (0.5452) -0.0341 (0.7905) 0.0825 (0.5204) 0.0779 (0.5440) -0.0854 (0.5056) -0.0636 (0.6206)
ONL+IS -0.1404 (0.2724) -0.0559 (0.6636) -0.2545 (0.0441) -0.0482 (0.7076) -0.2803 (0.0261) -0.0645 (0.6155) -0.1007 (0.4324) -0.0334 (0.7948)
EZ -0.0488 (0.7043) -0.0926 (0.4704) 0.0521 (0.6848) 0.1479 (0.2475) -0.0466 (0.7171) -0.0741 (0.5638) -0.0777 (0.5449) -0.0747 (0.5608)
OS -0.1058 (0.4093) 0.0478 (0.7097) -0.0853 (0.5063) 0.0486 (0.7054) 0.0889 (0.4885) 0.0977 (0.4463) 0.0400 (0.7554) 0.0952 (0.4578)
RPE -0.0816 (0.5249) 0.1665 (0.1922) -0.1138 (0.3743) -0.0633 (0.6219) -0.0143 (0.9117) 0.1426 (0.2650) -0.0028 (0.9826) 0.1983 (0.1192)
Fluid 0.0057 (0.9644) 0.1375 (0.2826) 0.1097 (0.3922) -0.0698 (0.5665) 0.1962 (0.1233) 0.2204(0.0827) 0.0952 (0.4580) -0.0928 (0.4694)
HRF -0.0037 (0.9770) 0.0765 (0.5510) -0.0286 (0.8239) -0.0945 (0.4615) 0.0787 (0.5398) 0.0783(0.5421) -0.0570 (0.6574) -0.1242 (0.3321)
Proliferative Diabetic Retinopathy (PDR)
RNFL 0.1601 (0.2825) 0.1992 (0.1794) 0.2971 (0.0426) 0.2557 (0.0827) 0.3684 (0.0108) 0.3445 (0.0177) 0.3349 (0.0214) 0.1851 (0.2130)
GCL+IPL -0.0184 (0.9024) -0.0757 (0.6129) -0.1019 (0.4953) -0.0155 (0.9177) -0.0330 (0.8259) -0.0793 (0.5963) -0.0604 (0.6867) -0.0456 (0.7607)
INL 0.1897 (0.2017) 0.0554 (0.7113) -0.0196 (0.8958) 0.1538 (0.3020) 0.0125 (0.9333) 0.2325 (0.1157) -0.0116 (0.9384) -0.0001 (0.9995)
OPL 0.1846 (0.2142) 0.2647 (0.0721) 0.1235 (0.4080) 0.2485 (0.0922) 0.3036 (0.0380) 0.2714 (0.0650) 0.2101 (0.1564) 0.2221 (0.1336)
ONL+IS 0.2279 (0.1234) 0.2712 (0.0652) 0.0472 (0.7528) 0.2578 (0.0802) 0.2041 (0.1688) 0.4315 (0.0025) 0.0320 (0.8311) 0.2863 (0.0511)
EZ 0.1297 (0.3850) 0.0653 (0.6628) 0.0527 (0.7250) 0.0458 (0.7596) 0.1073 (0.4730) 0.0588 (0.6946) 0.2042 (0.1686) 0.3334 (0.0220)
OS 0.0194 (0.8968) 0.0340 (0.8204) -0.0568 (0.7047) 0.0318 (0.8321) -0.0492 (0.7428) -0.0093 (0.9505) 0.1254 (0.4011) 0.1112 (0.4566)
RPE -0.0499 (0.7393) -0.0158 (0.9162) 0.0683 (0.6481) -0.0629 (0.6744) -0.0696 (0.6418) 0.1430 (0.3376) 0.2180 (0.1410) 0.2236 (0.1308)
Fluid 0.1768 (0.2346) 0.1607 (0.2806) 0.3300 (0.0235) 0.2877 (0.0499) 0.3010 (0.0398) 0.2914 (0.0469) 0.2255 (0.1275) 0.3372 (0.0204)
HRF 0.3477 (0.0166) 0.2440 (0.0984) 0.1824 (0.2198) 0.2670 (0.0696) 0.3777 (0.0089) 0.3964(0.0058) 0.2549 (0.0837) 0.1004 (0.5020)
TABLE IV: Partial correlation between retinal layer thickness and visual acuity in logMAR for NPDR and PDR. Each cell presents the Pearson correlation coefficient (p-value) adjusted by age, gender, and diabetes duration. Statistically significant values ($p<0.05$) are highlighted in bold.
Figure 10: Example of 62-year-old PDR male patient’s left eye image. The area enclosed by the red rectangle indicates the under-segmentation of HRF from the ground truth, where the model predictions have a better segmentation result.
V Conclusion
This study highlights the strengths and limitations of current segmentation models in analyzing DR-related structural changes. SwinUNETR and VM-UNet demonstrate strong performance, particularly in segmenting complex retinal layers and fluid regions, but challenges remain in HRF segmentation due to its small and widely distributed nature. The analysis of retinal layer thickness differences between NPDR and PDR reveals distinct structural alterations, with significant findings in RNFL, GCL+IPL, OS, fluid, and HRF distributions. The different responses of visual acuity within NPDR and PDR suggest a progression from adaptive retinal changes in NPDR to pathological neurodegeneration and edema-driven impairment in PDR, reinforcing the need for early intervention to prevent irreversible vision loss.
These results emphasize the potential for deep-learning models to improve disease monitoring and classification, and they provide valuable insights into the progressive nature of retinal changes in DR. However, addressing limitations such as manual labelling inconsistencies, cross-sectional study constraints, and limited NPDR severity stratification will be critical in refining segmentation algorithms and advancing the clinical utility of automated OCT analysis. Future efforts should focus on integrating longitudinal data, expanding datasets, and incorporating multi-modal imaging to enhance the predictive power of segmentation-based disease assessment tools.
VI Funding and acknowledgement
We would like to thank Roy Boustani, Talha Mohammed and Shuting Xing for the help of manual segmentation and corrections. This study was funded by the National Sciences and Engineering Research Council of Canada; Canadian Institutes of Health Research, Compute Canada, and Wake Forest University School of Medicine Startup Funding. The sponsor or funding organization had no role in the design or conduct of this research. The pipelines developed in this work will be available in our Cloud Engine Resource for Accelerated Medical Image Computing for Clinical Applications (CERAMICCA) platform (https://ceramicca.from-ca.com/).).
VII Declaration of competing interest
All co-authors declare no conflict of interest.